find the missing coordinate to complete the ordered pair solution to the given linear equations. 3x-2y=5 (a) (-3, ) (b) ( ,2)
Answer by stanbon(75887)

find the missing coordinate to complete the ordered pair solution to the given linear equations. 3x-2y=5
(a) (-3,y):: y = (3x-5)/2 ; y = (3*-3-5)/2 = (-9-5)/2 = -7
----------------------
(b) (x,2):: x = (5+2y)/3 ; x = (5+2*2)/3 = (5+4)/3 = 3
-----------------
Cheers,
Stan H.
-------------